OpenStudy (anonymous):

How do I find the ratio of the shortest side to the longest side if my numbers are 18 3/5, 9, 21. Then write the ratio in lowest terms with whole numbers in the numerator and denominator. (Show steps, please and thank you!)

OpenStudy (ash2326):

We are given \[18 \frac{3}{5}, 9, 21\] These are the three sides of a triangle. Let's first convert the mixed fraction into a decimal number \[18\frac{3}{5}= 18+\frac{3}{5}= 18 \times \frac{5}{5}+\frac{3}{5}=>\frac{90}{5}+\frac{3}{5}\] we get \[\frac{93}{5}=18.6\] Let's arrange in ascending order \[9, 18.6 , 21\] Smallest side=9 Longest side= 21 \[Ratio\ =\frac{9}{21}\] to convert this into lowest terms. Divide both numerator and denominator by 3 \[Ratio\ =\huge {\frac{\frac{9}{3}}{\frac{21}{3}}=\frac{3}{7}}\]

OpenStudy (anonymous):

Thank you so much Ash!

OpenStudy (ash2326):

@LynnMichael Welcome:)